OpenStudy (anonymous):

The perimeter of a rectangle is 34 inches. The length is 4 less than twice the width. What is the length of the rectangle. Write a system of equations for this problem. Which one of the equations below would be part of that system? DO NOT SOLVE

OpenStudy (saifoo.khan):

do you have any options?

OpenStudy (anonymous):

yea 2L + 2W = 4 2L + 2W = 30 L = 2W -4 L = 4W – 2

OpenStudy (saifoo.khan):

now the question is complete!! D/

OpenStudy (anonymous):

lozl sorry i shud have put it up b4 :)

OpenStudy (saifoo.khan):

No problem my friend! :D